# Primordial Black Holes and Scalar-induced Gravitational Waves in Sneutrino Hybrid Inflation

### v2

@@ Abstract
We investigate the possibility that primordial black holes (PBHs) can be formed from large curvature perturbations generated during the waterfall phase transition in a supersymmetric scenario where sneutrino is the inflaton in a hybrid inflationary framework. We obtain a spectral index ($n_{s}\simeq 0.966$), and a tensor-to-scalar ratio ($r\simeq 0.0056-10^{-11}$), consistent with the current Planck data satisfying PBH as dark matter (DM) and detectable Gravitational Wave (GW) signal. Our findings show that the mass of PBH and the peak in the GW spectrum is correlated with the right-handed (s)neutrino mass. We identify parameter space where PBHs can be the entire DM candidate of the universe (with mass $10^{-13}\,M_{\odot}$) or a fraction of it. This could be tested via second-order GW signals in future observatories, for instance, with amplitude $\Omega_{\rm GW}h^{2}$ $\sim 10^{-9}$ and peak frequency $f\sim 0.1$ Hz in LISA and $\Omega_{\rm GW}h^{2}\sim 10^{-11}$ and peak frequency of $\sim 10$ Hz in ET. We study two models of sneutrino inflation: Model11-1- 1 involves canonical sneutrino kinetic term which predicts the sub-Planckian mass parameter $M$, while the coupling between a gauge singlet and the waterfall field, $\beta$, needs to be quite large whereas, for the model22-2- 2 involving $\alpha-$attractor canonical sneutrino kinetic term, $\beta$ can take a natural value. Estimating explicitly, we show that both models have mild fine-tuning. We also derive an analytical expression for the power spectrum in terms of the microphysics parameters of the model like (s)neutrino mass, etc. that fits well with the numerical results. The typical reheat temperature for both the models is around $10^{7}-10^{8}$ GeV suitable for non-thermal leptogenesis.
@@ 1 Introduction
Following recent remarkable data from cosmic microwave background radiation (CMBR) acquired by the Planck satellite, cosmic inflation has become the paradigm for early universe cosmology. Moreover, due to the vacuum energy scale during inflation being approximately $10^{14}$ GeV (largest allowed energy scale by CMB) is close to the expected Grand Unification scale and seesaw scale, such measurements are a pathway to directly probe into ultra-violate particle physics.
@@
Alongside, the CMB scale and the physics during inflation, there has also been a great growing interest in the late stages of the inflationary paradigm, particularly the investigating scenarios where considerably large peaks in the amplitude of perturbations could be realized. This is because such peaks could lead to the formation of primordial black holes (PBHs) which has gained lots of interest after the observations of the supermassive  [1, 2] and stellar-mass black hole (BH) merger events detected via observing gravitational wave (GW) in LIGO-VIRGO-KAGRA [3, 4]. Such large primordial density fluctuation could generate a stochastic gravitational wave background (SGWB) when PBHs are generated via the collapse of high-dense regions [5, 6, 7]. Typically, this requires the inflaton to go through a very flat potential during inflation (see ref. [8] for a recent review on PBHs) and involves some higher degrees of fine-tuning; see ref.[9] for details. This was first studied decades ago in ref. [10] in the context of a hybrid inflation scenario [11, 12]. It was recently shown that hybrid inflation naturally reduces the fine-tuning involved in producing PBH as dark matter (DM) [13, 14].
@@
In this paper, we reconsider the possibility that the inflaton sector is responsible for generating the tiny neutrino masses required to explain several neutrino oscillation experiments. Heavy singlet neutrinos with scalar partners called sneutrinos with $10^{10}-10^{15}$ GeV, lies in the range where the inflaton mass may lie. In this paper, we discuss two supersymmetric scenarios where the lightest heavy singlet (s)neutrino drives inflation. As we will see, this scenario constrains in interesting ways many of the $18$ parameters of the minimal seesaw model for generating three non-zero light neutrino masses. This minimal (s)neutrino inflationary scenario can be rescued within the framework of hybrid inflation along with a successful neutrino oscillation data explanation. In terms of the early universe, this scenario provides a natural way to accommodate baryogenesis via leptogenesis due to sneutrino decay [15]. We will show that the mass of PBH and the peak in the GW spectrum is correlated with the right-handed (s)neutrino mass.
@@
As shown previously in the literature, the waterfall field can have a flat potential to generate large curvature perturbations at small scales, since the waterfall transition happens at the last stages of inflation [16, 17, 18]. These large density perturbations, lead to a larger amplitude of curvature perturbations and subsequently collapsing into a PBH. As shown in refs. [19, 20], this often leads to the overproduction of PBHs of astrophysical size. However, with a slightly modified waterfall field potential, this can be avoided and PBH abundance can be controlled with detectable scalar-induced Gravitational Wave (SIGW) signals [21]. Since there are several large uncertainties present in the critical process of PBH formation estimates, we study it in less detail. On the other hand, we study in detail the characterization of the SGWB induced at the second-order by the large curvature perturbations (at small length scales) during horizon re-entry in the radiation-dominated era [^1] See ref. [22, 23] for other ways to generate PBH during hybrid inflation scenarios..
@@ 2 Model11-1- 1: Toy Model Hybrid Inflation
Here, $\phi$ is the inflaton field, $\psi$ is the waterfall field, $\phi_{c}$ is the critical point below which the potential develops a tachyonic instability, forcing the field trajectories to reach one of the global minima, located at $\phi=0,\,\,\psi=\pm M$, $m_{1}$, $m_{2}$ are the dimensionful mass parameters and $\Lambda$ is the non-zero vacuum energy [^2] The minimum of the whole potential is $V_{\rm min}=\pm\Lambda(M/b)$. Now, $\Lambda$ is the inflation energy density scale because $H^{2}=\Lambda/3$ as mentioned before eq. 4.3, however, $b$ is not too small (in Table 1 and Fig. 3 we will see that $b\sim 10^{9}m_{\text{Pl}}$ and $M=0.1\,m_{\text{Pl}}$ (Table 1). This means that $V_{\text{min}}\sim\pm 10^{-22}m_{\text{Pl}}^{4}$. Although this is huge compared to the vacuum density of $\Lambda$CDM we can always adjust the unknown cosmological constant to cancel the vacuum density of the deeper minimum.. Usually, the hybrid inflation model predicts PBH overabundance [19]. We have added a linear term in the waterfall field that does not allow $\psi$ to relax exactly at zero but slightly displace depending upon the sign. of the co-efficient of the linear term $1/b$. In this way, one can control the peak of the power spectrum and avoid the PBH overproduction A schematic picture of the model is shown in fig. 1.
@@
The central measurements by Planck 2018 in the $\Lambda$CDM model are; $n_{s}=0.9647\pm 0.012$ and the tensor to scalar ratio $r=16\,\epsilon_{V}<0.035$ at $95\%$ C.L [25]. All the values are measured at the pivot scale, $k_{0}=0.05\,\text{Mpc}^{-1}$. The subscript 00 from here and onward indicates the value at the pivot scale. Along the valley, one obtains the slow roll parameters eq. 2.2 to be,
@@
where $\xi_{e}$ is the field value at the end of inflation which is fixed by the breakdown of the slow-roll approximation. The parameter set for different model variables we consider here is given in table 1.
@@ 3 Numerical Treatment of Scalar Perturbations
The Klien-Gordon classical background equations of motion in the number of e-fold times are given by [32],
@@
We, later on, use this relation to numerically evaluate the power spectrum.
@@ 4 Power Spectrum Analytical Expression
FIGURE 3: Power spectrum by solving the exact linear scalar perturbation equations (solid) eq. 3.7 and for the analytical expression (dashed) eq. 4.19 for the BP in table 1.
@@
Here, $\psi_{k}\equiv d\psi/d\chi=\psi_{0}\,e^{\chi_{k}}$. One can calculate $\chi_{k}$ by solving eq. 4.11 and eq. 4.9,
@@
$N_{1,\psi}=\dfrac{-M^{2}\,\sqrt{m_{1}}}{\sqrt{b^{-2}\,M^{4}\,m_{1}+32\,\phi_{c}\,\psi_{0}^{2}\,\chi_{2}}}\,\dfrac{1}{\exp{\left(\dfrac{8\,\left(N_{k}-N_{\text{e}}\right)^{2}}{M^{2}\,m_{1}\,\phi_{c}}-b_{1}^{-1}\,\left(N_{k}-N_{\text{e}}\right)\right)}}.$ (4.17)
@@
$P_{R}\simeq\dfrac{H^{2}}{4\,\pi^{2}}\left(\dfrac{m_{1}^{2}}{m_{\text{Pl}}^{4}}+\left(\dfrac{\mathbb{b}\,M^{2}\,\sqrt{m_{1}}\,e^{-\left(\dfrac{8\,m_{\text{Pl}}^{4}\left(N_{k}-N_{\text{e}}\right)^{2}}{M^{2}\,m_{1}\,\phi_{c}}-b^{-1}\,m_{\text{Pl}}\,\left(N_{k}-N_{\text{e}}\right)\right)}}{m_{\text{Pl}}^{3}\sqrt{b^{-2}\,M^{4}\,m_{1}+32\,\phi_{c}\,\psi_{0}^{2}\,\chi_{2}}}\right)^{2}\right).$ (4.19)
@@
Here, we introduce a fudge factor $\mathbb{b}=10^{-1}$ to keep the consistency of the previously used approximations and recover the Planckian units. In fig. 3, we present an exact power spectrum (numerically solved) along with the analytical expression given in eq. 4.19 for the potential eq. 2.1 and the BP in table 1. This clearly represents that the linear term helps to reduce the height of the power spectrum. The analytical expression is slightly off for the larger values of $b$ since estimations were made to obtain eq. 4.17. It is important to note that the analytical expression overestimates the power spectrum due to the fact that if we do not implement the assumption in eq. 4.11, the parameter $b^{-1}$ will evolve as a decreasing function due to the increasing behavior of the waterfall field. Therefore, the exponential factor would be suppressed. The width of the power spectrum is defined by the number of e-folds in the waterfall regime and hence is not affected by the underlying assumption.

### v4

@@ Abstract
We investigate the possibility that primordial black holes (PBHs) can be formed from large curvature perturbations generated during the waterfall phase transition in a supersymmetric scenario where sneutrino is the inflaton in a hybrid inflationary framework. We obtain a spectral index ($n_{s}\simeq 0.966$), and a tensor-to-scalar ratio ($r\simeq 0.0056-10^{-11}$), consistent with the current Planck data satisfying PBH as dark matter (DM) and detectable Gravitational Wave (GW) signal. Our findings show that the mass of PBH and the peak in the GW spectrum is correlated with the right-handed (s)neutrino mass. We identify parameter space where PBHs can be the entire DM candidate of the universe (with mass $10^{-13}\,M_{\odot}$) or a fraction of it. This can be tested in future observatories, for example, with amplitude $\Omega_{\rm GW}h^{2}$ $\sim 10^{-9}$ and peak frequency $f\sim 0.1$ Hz in LISA and $\Omega_{\rm GW}h^{2}\sim 10^{-11}$ and peak frequency of $\sim 10$ Hz in ET via second-order GW signals. We study two models of sneutrino inflation: Model11-1- 1 involves canonical sneutrino kinetic term which predicts the sub-Planckian mass parameter $M$, while the coupling between a gauge singlet and the waterfall field, $\beta$, needs to be quite large whereas, for the model22-2- 2 involving $\alpha-$attractor canonical sneutrino kinetic term, $\beta$ can take a natural value. Estimating explicitly, we show that both models have mild fine-tuning. We also derive an analytical expression for the power spectrum in terms of the microphysics parameters of the model like (s)neutrino mass, etc. that fits well with the numerical results. The typical reheat temperature for both the models is around $10^{7}-10^{8}$ GeV suitable for non-thermal leptogenesis.
@@ 1 Introduction
Following recent remarkable data from cosmic microwave background radiation (CMBR) acquired by the Planck satellite, cosmic inflation has become the paradigm for early universe cosmology. Moreover, due to the vacuum energy scale during inflation being approximately $10^{14}$ GeV (largest energy scale allowed by CMB) is close to the expected Grand Unification scale and seesaw scale, such measurements are a pathway to directly probe into ultra-violate particle physics.
@@
Alongside the CMB scale and the physics during inflation, there has also been a great growing interest in the late stages of the inflationary paradigm, particularly the investigating scenarios where considerably large peaks in the amplitude of perturbations could be realized. This is because such peaks could lead to the formation of primordial black holes (PBHs) which has gained a lot of interest after the observations of the supermassive  [1, 2] and stellar-mass black hole (BH) merger events detected via observing gravitational wave (GW) in LIGO-VIRGO-KAGRA [3, 4]. Such a large primordial density fluctuation could generate a stochastic gravitational wave background (SGWB) when PBHs are generated via the collapse of high-dense regions [5, 6, 7]. Typically, this requires the inflaton to go through a very flat potential during inflation (see ref. [8] for a recent review on PBHs) and involves some higher degrees of fine-tuning; see ref.[9] for details. This was first studied decades ago in ref. [10] in the context of a hybrid inflation scenario [11, 12]. It was recently shown that hybrid inflation naturally reduces the fine-tuning involved in producing PBH as dark matter (DM) [13, 14] [^1] Some other studies on PBHs include: Supermassive primordial black holes in the multiverse: for nano-Hertz gravitational wave and high-redshift JWST galaxies has been studied in [15]. The study of induced gravitational waves from flipped SU(5) superstring theory at nHz has been explored in [16]..
@@
In this paper, we reconsider the possibility that the inflaton sector is responsible for generating the tiny neutrino masses required to explain several neutrino oscillation experiments. Heavy singlet neutrinos with scalar partners called sneutrinos with $10^{10}-10^{15}$ GeV, lie in the range where the inflaton mass may lie. In this paper, we discuss two supersymmetric scenarios where the lightest heavy singlet (s)neutrino drives inflation. As we will see, this scenario constrains in interesting ways many of the $18$ parameters of the minimal seesaw model for generating three non-zero light neutrino masses. This minimal (s)neutrino inflationary scenario can be rescued within the framework of hybrid inflation along with a successful neutrino oscillation data explanation. In terms of the early universe, this scenario provides a natural way to accommodate baryogenesis via leptogenesis due to sneutrino decay [17]. We will show that the mass of PBH and the peak in the GW spectrum are correlated with the right-handed (s)neutrino mass given that (s)neutrino is responsible for driving the cosmic inflaton phase.
@@
As previously demonstrated in the literature, the waterfall transition happening in the last stages of inflation, can have a flat potential to generate large curvature perturbations at small scales [18, 19, 20]. These large density perturbations, lead to a larger amplitude of curvature perturbations and eventually collapses into a PBH. As demonstrated in refs. [21, 22], this usually leads to the PBH overproduction of astrophysical size. However, this can be circumvented and the PBH abundance modulated with observable Scalar Induced Gravitational Wave (SIGW) signals using a slightly modified waterfall field potential [23]. The PBH overproduction issue is addressed by introducing the linear term along with consistent CMB predictions. We present in our analysis the analytical expressions for the power spectrum (see section 4) in terms of the hybrid inflation microphysics parameters taking care of the PBH overproduction issue. Since the crucial process of PBH generation estimates is subject to multiple significant uncertainties, we study it in less detail. However, we study in detail the characterization of the SGWB induced at the second order by the large curvature perturbations (at small length scales) during horizon reentry in the radiation-dominated era [^2] See ref. [24, 25] for other ways to generate PBH during hybrid inflation scenarios. For grand unification in hybrid inflation, we refer the reader to [26]. A detailed analysis of the T and E models of hybrid inflation is discussed in [27]..
@@ 2 Model11-1- 1: Toy Model Hybrid Inflation
Here, $\phi$ is the inflaton field, $\psi$ is the waterfall field, $\phi_{c}$ is the critical point below which the potential develops a tachyonic instability, forcing the field trajectories to reach one of the global minima, located at $\phi=0,\,\,\psi=\pm M$, $m_{1}$, $m_{2}$ are the dimensionful mass parameters and $\Lambda$ is the non-zero vacuum energy [^3] The minimum of the whole potential is $V_{\rm min}=\pm\Lambda(M/b)$. Now, $\Lambda$ is the inflation energy density scale because $H^{2}=\Lambda/3$ as mentioned before eq. 4.3, however, $b$ is not too small (in Table 1 and Fig. 3 we will see that $b\sim 10^{9}m_{\text{Pl}}$ and $M=0.1\,m_{\text{Pl}}$ (Table 1). This means that $V_{\text{min}}\sim\pm 10^{-22}m_{\text{Pl}}^{4}$. Although this is huge compared to the vacuum density of $\Lambda$CDM we can always adjust the unknown cosmological constant to cancel the vacuum density of the deeper minimum.. Hybrid inflation models, where PBH formation was first presented rigorously [21], predict the overabundance of DM. We have added a linear term in the waterfall field that does not allow $\psi$ to relax exactly at zero but slightly displace depending upon the sign. of the co-efficient of the linear term $1/b$. In this way, one can control the peak of the power spectrum and avoid PBH overproduction. A schematic picture of the model is shown in fig. 1.
@@
The central measurements by Planck 2018 in the $\Lambda$CDM model are; $n_{s}=0.9647\pm 0.012$ and the tensor to scalar ratio $r=16\,\epsilon_{V}<0.035$ at $95\%$ C.L [29]. All the values are measured at the pivot scale, $k_{0}=0.05\,\text{Mpc}^{-1}$. The subscript 00 from here and onward indicates the value at the pivot scale. Using eq. 2.2, along the valley when $X=\phi$ one obtains the slow roll parameters to be,
@@
where $X_{e}$ is the field value at the end of inflation which is fixed by the breakdown of the slow-roll approximation. The parameter set for different model variables we consider here is given in table 1.
@@ 3 Numerical Treatment of Scalar Perturbations
The Klein-Gordon classical background equations of motion in the number of e-fold times are given by [36],
@@
Later we used this relation to numerically evaluate the power spectrum.
@@ 4 Power Spectrum Analytical Expression
Here, $\chi_{2}\equiv\ln\left(\sqrt{\phi_{c}/m_{1}}M/2\,\psi_{0}\right)$ and $\psi_{k}\equiv d\psi/d\chi=\psi_{0}\,e^{\chi_{k}}$. One can calculate $\chi_{k}$ by solving eq. 4.11 and eq. 4.9,
@@
$N_{1,\psi}=\dfrac{-M^{2}\,\sqrt{m_{1}}}{\sqrt{b^{-2}\,M^{4}\,m_{1}+32\,\phi_{c}\,\psi_{0}^{2}\,\chi_{2}}}\,\dfrac{1}{\exp{\left(\dfrac{8\,\phi_{c}\,\left(N_{k}-N_{\text{e}}\right)^{2}}{M^{4}\,m_{1}}-b_{1}^{-1}\,\left(N_{k}-N_{\text{e}}\right)\right)}}.$ (4.17)
@@
$P_{R}\simeq\dfrac{H^{2}}{4\,\pi^{2}}\left(\dfrac{m_{1}^{2}}{m_{\text{Pl}}^{4}}+\left(\dfrac{\mathbb{b}\,M^{2}\,\sqrt{m_{1}}\,e^{-\left(\dfrac{8\,m_{\text{Pl}}^{4}\,\phi_{c}\left(N_{k}-N_{\text{e}}\right)^{2}}{M^{4}\,m_{1}}-b^{-1}\,m_{\text{Pl}}\,\left(N_{k}-N_{\text{e}}\right)\right)}}{m_{\text{Pl}}^{3}\sqrt{b^{-2}\,M^{4}\,m_{1}+32\,\phi_{c}\,\psi_{0}^{2}\,\chi_{2}}}\right)^{2}\right).$ (4.19)
@@
Here, we introduce a fudge factor $\mathbb{b}=10^{-1}$ to keep the consistency of the previously used approximations and recover the Planckian units. In the left panel of fig. 3, we present an exact power spectrum (numerically solved) along with the analytical expression given in eq. 4.19 for the potential eq. 2.1 and the BP in table 1. This represents that the linear term helps to reduce the height of the power spectrum. The analytical expression is slightly off for the larger values of $b$ since estimations were made to obtain eq. 4.17. It is important to note that the analytical expression overestimates the power spectrum due to the fact that if we do not implement the assumption in eq. 4.11, the parameter $b^{-1}$ will evolve as a decreasing function due to the increasing behavior of the waterfall field. Therefore, the exponential factor would be suppressed. The width of the power spectrum is defined by the number of e-folds in the waterfall regime and hence is not affected by the underlying assumption.
As we know PBHs are formed in large over-dense regions, which means their formation rate can be significantly modified due to the non-Gaussian tails of the density perturbation distribution function as shown in several studies [38, 39, 40]. However, investigating non-Gaussian effects remains an open problem, receiving a huge interest in current research. In our present analysis, we do not attempt to include the effects of non-Gaussianity; rather we consider them as uncertainties. Below we present a discussion regarding the impact of non-Gaussianities. The first correction that arises if one goes beyond the Gaussian approximation is known as the scalar bi-spectrum $B_{R}$ which can be written as [41, 42]
$B_{R}\equiv\left\langle R_{\bm{k}_{1}}R_{\bm{k}_{2}}R_{\bm{k}_{3}}\right\rangle=(2\pi)^{3}\delta^{3}\left(\bm{k}_{1}+\bm{k}_{2}+\bm{k}_{3}\right)B_{R}\left(k_{1},k_{2},k_{3}\right),$ (4.20)
For detailed expressions of such bi-spectrum $B_{R}(k_{1},k_{2},k_{3})$ see [43, 44, 45]. Re-casting this in terms of the usual dimension-less $f_{\rm NL}$ parameter is expressed as [46, 41]
$f_{\text{NL}}(k_{1},k_{2},k_{3})=\frac{5}{6}\frac{B_{R}(k_{1},k_{2},k_{3})}{P_{R}(k_{1})P_{R}(k_{2})+P_{R}(k_{2})P_{R}(k_{3})+P_{R}(k_{3})P_{R}(k_{1})}.$ (4.21)
For the single field model, this implies the relation between spectral index and $n_{s}$ as $f_{\rm NL}|_{\rm squeezed}=5(1-n_{s})/12$ in the squeezed limit $k_{3}\to 0$ and thereby $k_{1}=k_{2}$ [47, 48]. The equilateral limit $k_{1}=k_{2}=k_{3}$ can be relevant for PBH formation for $P_{R}(k)$ which exhibits a peak, such that the PBH abundance gets approximately multiplied by [38, 49]
$\exp\left[-\frac{23\,\delta_{\rm c}^{3}}{P_{\zeta}(k_{\rm peak})}f_{\rm NL}(k_{\rm peak},k_{\rm peak},k_{\rm peak})\right].$ (4.22)
These considerations qualitatively suggest that unless $|f_{\rm NL}|\sim 1$, leading-order non-Gaussianities do not play a significant role in modifying PBH abundance. However, higher-order non-Gaussianities can be relevant with $R$ being expressed as a Gaussian fluctuation $R_{\rm G}$ in the local limit, such as those in ultra-slow roll inflationary models, see refs. [50, 40].
FIGURE 3: Left panel presents the power spectrum by solving the exact linear scalar perturbation equations (solid) eq. 3.7 and for the analytical expression (dashed) eq. 4.19 for the BP in table 1. The $f_{\text{NL}}$ parameter given in eq. 4.25 along with the power spectrum is presented in the right panel for the BP given in table 1 along with the variation of the coefficient of the linear term. The inset plot shows the peak value of the power spectrum on the log axis.
Here we present an estimate of $f_{\rm NL}$ using the $\delta N$ formalism following ref. [36]. The amplitude of scalar bi-spectrum is given by
$f_{\text{NL}}=-\dfrac{5}{6}\dfrac{\sum_{i,j}N_{,i}N_{,j}N_{,ij}}{\left(\sum_{i}N_{,i}^{2}\right)^{2}}.$ (4.23)
Here, $(i,j)$ refers to the derivative with respect to $(\phi,\psi)$. Considering phase 1, expanding the summation and using the chain rule, we arrive
$f_{\text{NL}}=-\dfrac{5}{6}\dfrac{N_{1,\phi}^{3}+N_{1,\psi}^{3}}{\left(N_{1,\phi}^{2}+N_{1,\psi}^{2}\right)^{2}}\,\dfrac{\partial N_{1,\psi}}{\partial N}.$ (4.24)
Using eqs. 4.18 and 4.17, it leads to the following analytical expression in $m_{\rm Pl}=1$ units,
$f_{\text{NL}}$ $\simeq-\dfrac{5}{6}\dfrac{\exp{\left(b^{-1}\,(N_{k}-N_{e})-\dfrac{8\,\phi_{c}\,(N_{k}-N_{e})^{2}}{M^{4}\,m_{1}}\right)}\left(M^{4}\,m_{1}\,b^{-1}-16\,\phi_{c}\,(N_{k}-N_{e})\right)}{m_{1}\left(M^{5}\exp{\left(2\,b^{-1}\,(N_{k}-N_{e})-\dfrac{16\,\phi_{c}\,(N_{k}-N_{e})^{2}}{M^{4}\,m_{1}}\right)}+M\,m_{1}\left(\dfrac{M^{4}m_{1}}{b^{2}}+32\,\phi_{2}\chi_{2}\psi_{0}^{2}\right)\right)^{2}}$ (4.25)
×$\times\left(M^{6}\exp{\left(3\,b^{-1}\,(N_{k}-N_{e})-\dfrac{24\,\phi_{c}\,(N_{k}-N_{e})^{2}}{M^{4}\,m_{1}}\right)}-m_{1}^{3/2}\left(\dfrac{M^{4}\,m_{1}}{b^{2}}+32\,\phi_{2}\chi_{2}\psi_{0}^{2}\right)^{3/2}\right).$
The resulting $f_{\text{NL}}$ along with the power spectrum is presented in the right panel of fig. 3 for the BP given in table 1. We have shown the variation of $b$ as indicated in the inset legends. At the CMB scale $f_{\text{NL}}$ is very well consistent with the experimental bound i.e. $f_{\text{NL}}(\text{CMB})\lesssim 4.2$ [51]. The maximum predicted value is $f_{\text{NL}}\lesssim\pm 1$ which corresponds to the small scales. Non-Gaussianities also somewhat modify the scalar-induced GWs as discussed in [52]. Again, unless $f_{\rm NL}\sim 1$ which may give an order unity correction to their frequency spectrum, see e.g. [52], such effects are expected to be small. However, fully non-Gaussian contributions arising from higher orders are still a topic of active research and debate, with new contributions being identified which have been previously overlooked in the literature [52, 53, 54]. A dedicated computation of the impact of such non-Gaussianities is beyond the scope of the current analysis and will be taken up in future studies.